OpenStudy (anonymous):

Can you find the leg of each isoseles right triangle when the hypotenuse is of the given measure of 5 square root of 6 and 3 must be the third number. Please help.

OpenStudy (anonymous):

\[\large 2a^2=c^2\]\[2a^2=(5\sqrt{6})^2\]\[2a^2=150\]\[a^2=75\]\[a=\sqrt{75}\]\[a=5\sqrt{3}\]

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

no not really